OpenStudy (anonymous):

A restaurant owner purchased x six-packs of soda. How many individual cans of soda did he purchase?

OpenStudy (anonymous):

Can someone help me out

OpenStudy (jdoe0001):

hmm how many "individual cans" in a six-pack anyway?

OpenStudy (anonymous):

6

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} six-pack&individuals\\ \hline\\ 1&6\cdot 1\\ x&6\cdot x \end{array}\)

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

so... see it ?

OpenStudy (jdoe0001):

you're not expected to have an absolute quantity, just an expression

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} six-pack&individuals\\ \hline\\ 1&6\cdot 1\implies 6\\ x&6\cdot x\implies {\color{red}{ 6x}} \end{array}\)

OpenStudy (anonymous):

so I write that for the answer

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

ok